OpenStudy (anonymous):

find f'x for f(x)= (3x-sin^2(4x))^1/2

OpenStudy (anonymous):

Alright...you have: \[f(x)=(3x-\sin^2(4x))^{\frac{1}{2}}=\sqrt{3x-\sin^2(4x)}\] You are going to be using the chain rule a bit here; i'll simplify a little if you don't mind: \[\eqalign{ f(x)&=\sqrt{3x-\sin^2(4x)} \\ &=\sqrt{3x-[\sin(4x)]^2} \\ &=\sqrt{3x-[g(x)]^2} \\ &=\sqrt{3x-h(x)} \\ &=\sqrt{j(x)} \\ }\] So then we can expand a little: \[\eqalign{ f'(x)&=\frac{d}{dx}\sqrt{j(x)} \\ &=\frac{1}{2\sqrt{j(x)}}\times j'(x) \\ &=\frac{1}{2\sqrt{j(x)}}\times[3-h'(x)] \\ &=\frac{1}{2\sqrt{j(x)}}\times[3-2g(x)\times g'(x)] \\ &=\frac{1}{2\sqrt{j(x)}}\times[3-2g(x)\times 4\sin(x)] \\ &=\frac{3-2g(x)\times 4\sin(x)}{2\sqrt{j(x)}} \\ &=\frac{3-8\sin(4x)\sin(x)}{2\sqrt{3x-\sin^2(4x)}} }\]

OpenStudy (anonymous):

hmmm Im a bit confused by the first part